# An Improved Two-Step Attack on Kyber

### v2

@@ 4. Experiments
The proposed attack is applied to Kyber512, and it is noteworthy that our method is also applicable to other security levels due to their use of the same $basemul$. We run the reference implementation from the pqm4 with -o compilation optimization on the STM32F407-DISCOVERY, an ARM Cortex-M4 microcontroller, at $48$MHz. Power traces were captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a sampling rate set to $500$ MSa/s. Based on the analysis in Section 3, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
@@
Figure 3. (a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$. Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
@@
As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focus on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions for 128 times. In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls.
To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The former is shown in Fig. 2 (a) and (b) of Section 3.1 and the modified model can reinforce the PCC value of the correct key. The latter shown in Fig. 2 (c) and (d) is to demonstrate its superiority in identifying the complementary false positive.
When the number of random ciphertexts $D$ is set to $15$, the relationship of the maximum ||||PCC|||| to the guess keys is shown in Fig. 2 (c). We can directly pick out the correct key $2280$ and its complementary false positive $1049$ (equal to $3329-2280$), which is a general phenomenon. In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated. To escape such false positive, we have modified the leakage model as described in Section 3.1. The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig. 2 (d). It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one. Consequently, to avoid the false positive one is selected, we should focus solely on the positive peak value of PCC. We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one. The intermediate values for the two guessed key coefficients are summarized below:
(4) $r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$
(5) $r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$
@@
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Fig. 3 (a) shows the points of the maximum ||||PCC|||| results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close. The phenomenon can lead to an inability to distinguish the correct coefficient. This is another kind of false positives. If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure. Therefore, we use the method mentioned in Section 3.2 to compensate for such case. First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time of this computation is negligible when compared to the complete PCC computing. Fig. 3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$. It shows that Kendall’s tau expands the differences between correct coefficient and false positives. As the number of power traces increases, the value of correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly. We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result. In another word, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces. Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 3 (c) and (d). It can be seen that the correct key $564$ can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we may not be able to figure it out any more.
Figure 4. The success rate and efficiency of our attack in twenty experiments.
@@
We conduct the trial-and-error lattice attack using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz. The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$. We collect the CPA results of the first sixty $basemuls$ calls. Fig. 4 shows the relationships between the success rate/time and the number of power traces represented in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful. Those cases with traces number smaller than $15$ can be compensated to some degree by running more trials of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the traces number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack are executed in serial. Therefore, we can summarize that when the success rate is large enough (close/equal to $1$), we can only cost at most $15$ power traces for a guess/call and about $9$ minutes to recover the full key on Kyber512.
Table 1. SCAs on Kyber to Recover a 128-Value Group of $\mathbf{s}$
@@
(xu2021magnifying, ) Pro SPA $\times$ INTT, ME 8$\sim$960 128
(mu2022voltage, ) Pro TA $\times$ MR 11 128
(sim2022chosen, ) Non-Pro SPA $\times$ MR 6$\sim$12 128
(yang2023chosen, ) Non-Pro CPA $\times$ CM 20$\sim$30 128
(kuo2023lattice, ) Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
@@
Table 1 shows the comparison results of SCAs on Kyber to recover a 128-value group of $\mathbf{s}$. The attack types include the profiled SPA (xu2021magnifying, ), profiled TA (mu2022voltage, ), non-profiled SPA (sim2022chosen, ), and non-profiled CPA (kuo2023lattice, ; yang2023chosen, ). Our method belongs to the non-profiled CPA, which does not cost extra training step. Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme (ravi2024side, ). The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme. For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work (yang2023chosen, ) which requires at least $20\times 128=2560$ power traces. If we slightly increase the traces number of the guess, the required time can be reduced to approximate $5$ minutes, much faster than the $20$ minutes recorded in (kuo2023lattice, ).
5. Conclusion
In this brief, we propose an efficient two-step attack including an enhanced CPA and a trial-and-error lattice attack for Kyber. In traditional CPA, false positives are the main factor affecting coefficients of secret key extraction in experiments (yang2023chosen, )and (kuo2023lattice, ). However, by employing the comprehensive CPA step we proposed, we can cleverly avoid false positives while reducing the number of required power traces in the practical attack. In the lattice step, we construct a trial-and-error algorithm and dynamically compute the lattice attack to reduce the power traces and time. Experimental results show that our proposed attack can accurately recover the full secret key of Kyber512 in about $9$ minutes with about 15 power traces of a guess on a machine with sixteen threads. Because the $basemul$ function called by different security levels of Kyber is the same, our work can be directly applied to the other parameters. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography.

### v3

@@ IV Experiments
We run the reference implementations of Kyber-512, -768, and -1024 from the pqm4 library [11] with -o compilation optimization on a STM32F407-DISCOVERY target board at $48$ MHz. The power traces are captured using a PicoScope 3043D oscilloscope and a CYBERTEK EM5030-2 electromagnetic probe, with a $500$ MSa/s sampling rate. Based on the analysis in Section III, we have chosen $fqmul(k,u_{1})$ as the intermediate value of our attack.
@@
Figure 3: (a) and (b) present the PCC and Ken results, respectively, for a another $basemul$ call when $D=15$. Similarly, (c) and (d) show the PCC and Ken results for the same $basemul$ call when $D=11$.
@@
As aforementioned, the CPA is divided into two stages: the capturing stage and the modeling computation stage. Our optimizations are mainly focused on the latter stage, including modifying the leakage model and adding an optional Kendall’s tau computing following the PCC computing.
For the capturing stage, we carry out the CPA at the NTT operation which calls $basemul$ functions 128 times. In our CPA experiments, we capture random ciphertexts and tremendous data of electromagnetic radiations during those calls. To show the effectiveness of the modified leakage model in the modeling computation stage, we have implemented two experiments. The former is shown in Fig. 2 (a) and (b) of Section III-A and the modified model can reinforce the PCC value of the correct key. The latter shown in Fig. 2 (c) and (d) is to demonstrate its superiority in identifying the complementary false positive.
When the number of random ciphertexts $D$ is set to $15$, the relationship of the maximum ||||PCC|||| to the guess keys is shown in Fig. 2 (c). We can directly pick out the correct key $2280$ and its complementary false positive $1049$ (equal to $3329-2280$), which is a general phenomenon. In addition, as the number of power traces decreases, the attacker would run into some problems because both the correct coefficient $s_{0}$ and its complementary value $q-s_{0}$ would have high and close PCC values, as the HW$(s_{0})$ and HW$(q-s_{0})$ are highly correlated. To escape such false positive, we have modified the leakage model as described in Section III-A. The correspondence between the PCC values of the correct key coefficient and the false positive one is shown in Fig. 2 (d). It is observed that the positive peak value corresponds to the correct key coefficient and the negative peak value of PCC corresponds to the false positive one. Consequently, to avoid the false positive being selected, we should focus solely on the positive peak value of PCC. We briefly explain this phenomenon as follows: Let $s_{0}$ be the correct one, and $s_{0}^{\prime}=3329-s_{0}$ be the false positive one. The intermediate values for the two guessed key coefficients are summarized below:
$r_{1}=\text{fqmul}(s_{0},u_{1})=(s_{0}\times u_{1})\,mod\,q,$ (5)
$r_{1}^{\prime}=\text{fqmul}(s_{0}^{\prime},u_{1})$ $=((3329-s_{0})\times u_{1})\,mod\,q$ (6)
@@
Figure 4: The success rate and efficiency of our attack in twenty experiments.
To further improve the accuracy in the CPA step, we adopt the Kendall’s tau to the picked candidates after the PCC process. Fig. 3 (a) shows the points of the maximum ||||PCC|||| results to the guess keys during another $basemul$ call, where $D=15$ and the number of candidates after PCC is equal to 4. It can be seen that the correlations of four coefficients $564$, $1687$, $2010$, and $2264$ stand out from the crowd of guessed coefficients when setting the threshold to $0.930$ and their correlation values are very close. The phenomenon can lead to an inability to distinguish the correct one. This is another kind of false positive. If the attacker just picks one of the four candidates at random, the accuracy is only about $25$%, which is still regarded as a failure. Therefore, we use the method mentioned in Section III-B to compensate for such a case.
First of all, we compute the Kendall coefficient matrix of these four candidate coefficients. It should be noted that the time of this computation is negligible when compared to the complete PCC computing. Fig. 3 (b) shows the values of the four maximum ||||Kendall’s tau|||| when $D=15$. It shows that Kendall’s tau expands the differences between the correct coefficient and false positives. As the number of power traces increases, the value of the correct coefficient tends to be more stable while the ||||Kendall’s tau|||| of false positives decrease quickly. We can easily distinguish the correct key $564$ when $D=15$ and output it as the final result. In other words, an attacker can take the correct coefficient $564$ for the $basemul$ call using at most 15 power traces. Similarly, we reduce the number of power traces to $11$ step by step and implement the PCC and Kendall’s tau computations. The corresponding results are shown in Fig. 3 (c) and (d). It can be seen that the correct key $564$ can almost not be recognized. We can pick out the final correct candidate from those four candidates according to the results of the summations of their two kinds of corrections. Note that if we continue reducing the number of power traces, we may not be able to figure it out anymore.
@@
We conduct the trial-and-error lattice attack for a 128-value group of $\mathbf{s}$ (hereinafter referred to as SK-128) using $16$ threads on a Xeon Gold 6342 @ $2.8$ GHz. The parameter of the $block$ size in the BKZ reduction is set to $50$ and the value of $max\_loops$ to $8$. We collect the CPA results of the first sixty $basemuls$ calls. Fig. 4 shows the relationships between the success rate/time and the number of power traces represented in the blue curve and groups of histograms, respectively. The success rate is evaluated by repeatedly conducting the proposed attack many times and counting the success times. It can be seen that when the number of power traces is no less than $15$, all the experimental attacks are successful. Those cases with trace numbers smaller than $15$ can be compensated to some degree by running more trials of the lattice attack. Meanwhile, according to the decomposed time histograms in Fig. 4, the time of the proposed lattice attack heavily relies on the number of power traces, while the CPA is almost unchanged although the trace number is doubled. The main reason could be the different utilization ratios of the multi-threading server. Note that the CPA computations are independently executed with the different power traces while the trial computations of the lattice attack of SK-128 are executed in serial.
Table I presents the time of implementing the proposed attacks on Kyber at different security levels, using 15 power traces across twenty experiments. Kyber-512, -768, and -1024 have 4, 6, and 8 SK-128s, respectively, with each SK-128 being independent of the others. This allows us to recover the different SK-128s in parallel. The attack method we propose can recover the full secret key on different security levels of Kyber, with the success rate remaining consistent. The reason for the slight increase in time taken at the high security level is the high parallelism of Step1 and the duration of Step2 fluctuates around 5 minutes. It is worth noting that reducing the parallelism will increase the time required for Kyber-768 and Kyber-1024, but the success rate remains unaffected.
TABLE I: Time Consumed to the Proposed Attacks on Kyber by Using 15 Power Traces in Twenty Experiments ($n=60$)
Security Level Step1 Step2 Total Time ($min$)
Time / Thread Time / Thread
SK-128 4.76 / $16\times 1$ 4.67 / $1\times 1$ 9.43
Kyber-512 5.01 / $16\times 4$ 4.66 / $1\times 4$ 9.67
Kyber-768 5.14 / $16\times 6$ 5.00 / $1\times 6$ 10.14
Kyber-1024 5.30 / $16\times 8$ 5.33 / $1\times 8$ 10.63
TABLE II: SCAs on Kyber to Recover SK-128
@@
[13] Pro SPA $\times$ INTT, ME 8$\sim$960 128
[14] Pro TA $\times$ MR 11 128
[15] Non-Pro SPA $\times$ MR 6$\sim$12 128
[7] Non-Pro CPA $\times$ CM 20$\sim$30 128
[8] Non-Pro CPA $\checkmark$ CM 800 50$\sim$87
@@
Table II shows the comparison results of SCAs on Kyber to recover SK-128. The attack types include the profiled SPA [13], profiled TA [14], non-profiled SPA [15], and non-profiled CPA [7, 8]. Our method belongs to the non-profiled CPA, which does not cost extra training step. Moreover, we only adopt random ciphertext, and thus we do not need extra effort like the oracle-based assistance to apply to the KEM scheme [2]. The restriction ciphertext CPA requires oracle-based assistance to be applied to the KEM scheme. For the performance, as shown in the table, the total number of required power traces of ours is $[11,15]\times 60=[660,900]$, much smaller than the state-of-the-art work [7] which requires at least $20\times 128=2560$ power traces. If we slightly increase the traces number of the guess, the required time can be reduced to approximately $5$ minutes, much faster than the $20$ minutes recorded in [8].
V Conclusion
In this brief, we propose an efficient two-step attack for Kyber, combining an enhanced CPA and a trial-and-error lattice attack. Experimental results show that the proposed attack can accurately recover the full secret key of Kyber-512, -768, and -1024 in about $9\sim 10$ minutes with about 15 power traces. Moreover, the core idea of the proposed attack is a general methodology and can be easily extended to other lattice-based cryptography. It is worth noting that our proposed attack method can be applied to implementations on hardware like FPGAs by slightly modifying the leakage model. To mitigate these attacks, algorithm-level protections such as shuffling [16] and masking [17] can be employed to protect the secret key, though these methods may involve trade-offs in efficiency and effectiveness.